# PointDreamer: Zero-shot 3D Textured Mesh Reconstruction from Colored Point Cloud by 2D Inpainting

### v1

@@ 4.2. Baselines
c—c—cccc Dataset Method PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
GSO SPR 27.11 0.907 0.1200 27.43
NKSR 24.65 0.900 0.1184 36.51
T.F. 25.53 0.894 0.1276 46.44
Ours 27.24 0.923 0.0827 9.32
OmniObject3D SPR
30.20 0.927 0.0998 40.13
NKSR 26.77 0.919 0.1048 50.14
T.F. 28.08 0.914 0.1121 65.61
Ours 29.78
0.941 0.0683 18.23
4.3. Main Results
4.3.1. Qualitative comparisons
Since DHSP3D is much slower than other methods (about 4-6 hours per shape compared to ¡ 2 min), we randomly select two objects from GSO dataset for efficiency when comparing our method against DHSP3D; see Figure 8. For other methods (i.e., SPR, NKSR & texture field), we present the visual comparisons on each dataset separately; see Figures 56 and 7.
Figures 57 show that our method, though zero-shot, achieves much clearer and more realistic textures than baselines, thanks to the strong 2D diffusion prior. Besides, SPR and NKSR, which both rely on per-point normals as input, sometimes yield redundant geometry mainly due to the wrongly-estimated normal direction. The texture field baseline, which is trained on ShapeNetCoreV2, performs much worse on the newly-seen dataset of GSO and Omniobject3d, indicating a lack of generalization ability.
In Figure 8, the red cup in the first row shows that DHSP3D does not support arbitrary topology. This is because it optimizes meshCNN and 2D XYZ map from the convex hull of the input point cloud, enforcing the output mesh to adhere to the hull’s topology. Besides, DHSP3D suffers from less realistic textures with jagged or unclear edges, since its purely self-prior property only enables it to predict point colors considering corresponding neighbors, without the ability to dream the unseen like our PointDreamer.
4.3.2. Quantitative comparisons
We summarize the quantitative results in Tables 4.2 and 4.2. We can see that our method outperforms baselines on most metrics and datasets, especially on the two perceptual metrics FID and LPIPS with a significant margin. However, regarding PSNR, which prioritizes pixel-level accuracy and thus may differ from human perception of quality, our method is sometimes outperformed by SPR. We assume that SPR predicts a point’s color by somehow fusing nearby points’ colors, leading to blurry textures but safely avoiding extreme pixel-level errors. On the contrary, our method dreams the colors by its diffusion prior. While achieving visually better results, a few pixels with extremely high errors may lower the overall PSNR score.
4.4. Impact of Degraded Input Quality
4.4.1. Anti-Noise Ability Analysis
We compare the anti-noise abilities of our method and baseline methods by applying them on point clouds with manually added Gaussian noise (standard deviation = 0.005) from the chair category of ShapeNetCoreV2 dataset. Table 4.5 shows the results, where our method, even with noisy input, outperforms baselines with clean input, indicating a strong anti-noise ability. Visual comparisons can be found in our supplementary file.
4.4.2. Sparsity Test
To evaluate our method’s ability to deal with sparse input, we decrease the input point number from 30k to 20k gradually, and present the results in Table 4.5. We can see only a small performance drop, and even with only 20k points as input, our method outperforms baseline methods with 30k points as input; see results in Table 4.2 “Chairs” category. Again, visual comparisons can be found in our supplementary file.
4.5. Method Component Analysis
We here compare different implementations for sub-modules (geometry extraction, inpainting and unprojection module) of our method by replacing one module at a time while keeping others unchanged. All experiments in this subsection are conducted on the full test set of the chair category of ShapeNetCoreV2 dataset with Gaussian noise (standard deviation = 0.005).
Table 3. Comparisons on anti-noise ability on chair category of ShapeNetCoreV2 dataset. “Noisy” denotes adding Gaussian noise with a standard deviation of 0.005 to the input point cloud.
@@ 4.5. Method Component Analysis
c—cc—cc—cc—cc
Method PSNR ↑   SSIM ↑   LPIPS ↓   FID ↓
Clean Noisy Clean Noisy Clean Noisy Clean Noisy
SPR 24.10 19.64 0.9311 0.8840 0.0923 0.1702 12.03 65.25
NKSR 23.05 22.79 0.9307 0.9288 0.0964 0.1017 29.11 44.02
T.F. 26.12 26.01 0.9467 0.9436 0.0664 0.0711 6.83 8.21
Ours 26.29 26.26 0.9524 0.9516 0.0570 0.0574 4.90 4.93
Table 4. Comparisons on our method’s performance with increased input point cloud sparsity.
{NiceTabular}
c—cccc Point Number PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
30k 26.2910 0.9524 0.0570 4.9048
25k 26.2761 0.9517 0.0580 5.2412
20k 26.2553 0.9509 0.0594 5.6960
4.5.1. Geometry Extraction
In addition to POCO, we further explore extracting geometry by SPR and depth inpainting to investigate how the geometry extraction module influence the final reconstructed textured mesh. Depth inpainting refers to inpainting projected sparse 2D depth maps instead of RGB images, and then reconstructing geometry from the inpainted dense depth maps by depth fusion (Griegler, 2017). Please refer to our supplementary file for more details and the quantitative results. We present the visual comparisons in Figure 9. We can see that, compared to POCO, both SPR and depth inpainting suffer from noisy geometry, resulting in noisy textures. Clearly, a higher-performance geometry extraction module contributes to a more refined reconstructed textured mesh.
Figure 9. Geometry extraction module replacement results comparing our POCO against SPR and Depth Inpainting.
4.5.2. 2D Inpainting
We compare different inpainting modules in our pipeline, including nearest interpolation, linear interpolation, DiffPIR (Zhu et al., 2023) and our adopted DDNM, where DiffPIR is another diffusion-based 2D image restoration method with inpainting ability. We provide the visual comparisons in Figure 10 and quantitative results in our supplementary file. As can be seen, the other three methods, though perform overall reasonably, fail to produce as clear textures as DDNM, indicating the importance of a strong inpainting module to the reconstruction performance.
Figure 10. Inpainting module replacement results comparing our adopted DDNM against nearest interpolation, linear interpolation, and DiffPIR.
4.5.3. Unprojection
We present the comparison results of different unprojection strategies in Figure  11 and Table 4.5.3.
  • “Opt. Scratch” denotes optimizing the texture atlas from scratch (Tang et al., 2023a);
  • “Naive” selects the best views by direction priority;
  • “NBF” selects the best views by our proposed Non-Border-First strategy;
  • “Opt. Naive” further optimizes the obtained atlas of “Naive”;
  • “Opt. NBF” further optimizes the obtained atlas of “NBF” while considering only loss in non-border areas.
The first row of Figure 11 shows an example, where optimizing from scratch fails on areas unseen from any view due to occlusion. Both rows show that the naive unprojection yields the above-mentioned artifacts from border areas, which can be slightly reduced by further optimization. Only our NBF and Opt. NBF nearly eliminate such artifacts. Table 4.5.3 also shows that our proposed NBF and Opt. NBF outperform other unprojection strategies.
Figure 11. Qualitative results of replacing our unprojection strategy with others. “Opt.” denotes “Optimize”.
Table 5. Quantitative results of replacing our NBF unprojection strategy with others. “Opt.” denotes “Optimize”.
@@ 4.5.3. Unprojection
Opt. Scratch (Tang et al., 2023a) 26.1939 0.9473 0.0592 5.2619
Naive 26.2225 0.9500 0.0587 5.1068
Opt. Naive 26.2477 0.9511 0.0580 4.9389
NBF 26.2729 0.9512 0.0579 5.0309
Opt. NBF 26.2565
0.9516 0.0574 4.9326
5. Conclusion, Limitations and Future Work
Conclusion. We propose PointDreamer, a novel zero-shot framework for textured mesh reconstruction from colored point cloud. PointDreamer predicts colors by 2D inpainting to utilize the strong diffusion prior and achieves SoTA performance. We also propose a novel “Non-Border-First” strategy to unproject the colors of predicted 2D images back to 3D space.
Limitation and Future Work. The main shortcoming of our proposed method is its relatively low speed. We employed one NVIDIA GeForce GTX 3090 GPU and it takes about 100s for our method to reconstruct a textured mesh, compared to NKSR’s 0.64 s. It takes about 72 s for DDNM (Wang et al., 2022b) to inpaint our 8 multi-view images, which takes most of the time. Replacing DDNM with a faster inpainting module may increase our reconstruction speed. In the future, we may seek to address this issue, and further explore 3D-aware inpainting and relighting-supported texture reconstruction.
Appendix A Appendix: Implementation Details
A.1. Texture Field Baseline
A.1.1. Network architecture
Inspired by works (Gupta et al., 2023; Shen et al., 2021) that represent 3D information by a feature tri-plane, we adopt the network architecture of the open-source Convolutional Occupancy Network (Peng et al., 2020), which also follows a tri-plane representation. Specifically, we modify its one-dimensional output head for occupancy prediction to three-dimensional for RGB color prediction.
A.1.2. Training
We follow 3DGen (Gupta et al., 2023) to train the network by the per-3D-point MSE loss on predicted and GT colors of sampled 3D points. Additionally, we also tried to employ the per-2D-pixel MSE loss of 2D images rendered from GT meshes and predicted meshes by differentiable rendering, but the experimental results are worse, as shown in Figure 12. Therefore, we adopt the per-3D-point MSE loss in our manuscript. We train the texture field network on a single NVIDIA GeForce GTX 3090 GPU with a batch size of 24 and a learning rate of 0.0001 for 2592, 024 iterations (about 356 epochs).
A.1.3. Inference.
During inference, we first use POCO (Boulch and Marlet, 2022) to predict an untextured mesh from the input point cloud, and apply UV mapping to it by Xatlas (Young, 2022), which produces the 3D positions of each valid pixel in the texture atlas. We query the color of each of these 3D positions by our trained texture field network, to inpaint the texture atlas, so as to obtain the final textured mesh.
Figure 12. Visual comparisons of the Texture Field baseline trained with different losses. “2D per-pixel” denotes rendering the generated textured mesh to multi-view 2D images, and then calculating the MSE loss between the rendered and GT images. “3D per-point” denotes calculating MSE loss between the predicted and GT colors of sampled 3D points.
A.2. Depth Inpainting for Geometry Extraction
@@
With 2D inpainting adopted as the key for our texture reconstruction, we can also introduce it to geometry reconstruction, by inpainting 2D depth maps instead of RGB images. Specifically, we follow the following steps to reconstruct untextured meshes from the input point cloud, as shown in Figure 13:
Figure 13. Pipeline of extracting geometry by depth inpainting.
  1. (1)
@@
  4. (4)
    We produce an untextured mesh by depth fusion based on the inpainted dense depth maps, and conduct mesh simplification (Garland and Heckbert, 1997) and Taubin Smooth (Taubin, 1995) to it as post-processing, to get the final untextured mesh.
Appendix B Appendix: More Experimental Results
B.1. Effect of Different $K$ Values (Number of Viewpoints for Projection and Inpainting)
To investigate the effect of different numbers of viewpoints (denoted as $K$) for projection and inpainting, we conduct experiments on the motorbike category or ShapeNetCoreV2 dataset by using different $K$ values including 6, 8, and 20. Table B.1 shows the distribution of cameras for each setting, together with the quantitative results. We can see that more views contribute to a slightly higher reconstruction quality. Visual comparisons in Figure 14 also show that, an insufficient number of views would produce artifacts in invisible or occluded areas, thus impacting the performance.
@@ B.1. Effect of Different $K$ Values (Number of Viewpoints for Projection and Inpainting)
Table 6. Quantitative results of using different $K$ values (numbers of viewpoints for projection and inpainting) on the motorbike category of ShapeNetCoreV2 dataset. More views contribute to a slightly higher reconstruction quality.

### v2

@@ 4.4 Multi-view Consistency
cc—cccc
Input Method PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
Noisy SPR 19.64 0.884 0.170 65.25
Noisy NKSR 22.79 0.929 0.102 44.02
Noisy T.F. 26.01 0.944 0.071 8.21
Noisy Ours 26.26 0.952 0.057 4.93
Clean T.F. 26.12 0.947 0.066 6.83
TABLE III: Performance comparison on ShapeNetCoreV2 Chairs with increased input point cloud sparsity.
@@
cc—cccc
Point Number Method PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
30k Ours 26.29 0.9524 0.057 4.90
25k Ours 26.28 0.9517 0.058 5.24
20k Ours 26.26 0.9509 0.059 5.70
10k Ours 26.17 0.9477 0.064 7.79
30k T.F. 26.12 0.9467 0.066 6.83
TABLE IV: Quantitative results of replacing our NBF unprojection strategy with others. “Opt.” denotes “Optimize”.
@@
Opt. Scratch 26.19 0.9473 0.0592 5.26
Naive 26.22 0.9500 0.0587 5.11
Opt. Naive 26.25 0.9511 0.0580 4.94
NBF 26.27 0.9512 0.0579 5.03
Opt. NBF 26.26 0.9516 0.0574 4.93
Figure 5: Our PointDreamer’s robustness against noisy and sparse input.
Figure 6: Our PointDreamer’s unique completion ability. When the input point cloud is incomplete e.g. scanned from occluded objects as shown in (a), only our PointDreamer can plausibly complete the missing part with high-quality texture, as demonstrated in (c). This is benefited from the strong inpainting power of the adopted 2D diffusion model, as illustrated in (b).
Figure 7: Sub-Module Replacement Results: (a) Geometry module. (b) Inpainting module. (c) Unprojection module.
Figure 8: Examples of adopting our proposed NBF strategy on other methods (InstantMesh [51]). (a) The original mesh generated by InstantMesh looks blurry. (b) Directly unprojecting intermediate multiview images to the mesh alleviates the blurriness, but leads to artifacts at the border of the eye as marked in a red box. (c) Using our NBF unprojection strategy avoids such artifacts. (d-g) More examples.
4.5 Robustness against Degraded Input Quality
4.5.1 Anti-Noise Ability Analysis
We compare our method and baseline methods by applying them on point clouds with manually added Gaussian noise (standard deviation = 0.005) from the chair category of ShapeNetCoreV2 dataset. Table 4.4 shows the results, where our method, even with noisy input, outperforms baselines with clean input. Figure 5 (b) also shows our PointDreamer’s strong anti-noise ability. While baseline methods exhibit significant performance drops given noisy input, our PointDreamer’s performance degradation is minimal: even barely noticeable by human perception.
4.5.2 Sparsity Test
To evaluate our method’s ability to deal with sparse input, we decrease the input point number from 30k to 10k gradually, and present the results in Table 4.4. We can see only a small performance drop, and even with only 20k points as input, our method outperforms baseline methods with 30k points as input. When the input point number is decreased to 10k, our PointDreamer achieves higher PSNR and SSIM metrics compared to the Texture Field baseline with 30k points. We also provide visual comparisons in Figure 5 (c), where our method significantly outperforms baseline methods with only 10k points as input.
4.5.3 Texture Completion Ability
Thanks to the inpainting power of the adopted 2D diffusion model, our PointDreamer has a unique ability to complete textures from incomplete scans. Figure 6 shows an example, where the point cloud of the table is incomplete since a small area is occluded by the cup during scanning. As shown in Figure 6 (a), the tablecloth is partially missing, revealing the black underside of the table below. For geometry reconstruction, such incompletion is less challenging: both our method and baseline methods manage to fill in this hole, as shown in Figure 6 (c). If facing more challenging cases, we can also add a point cloud completion module [52] before mesh reconstruction. However, texture completion is less explored, and baseline methods cannot effectively deal with it. Luckily, as shown in Figure 6 (b), our diffusion-based inpainting module addresses this issue. When projecting the incomplete 3D point cloud into a 2D sparse image from the top view, the occluded part is empty to be inpainted, thanks to our hidden point removal operation. The strong diffusion prior handles this empty area well and the inpainted dense image looks plausible. As a result, as shown in Figure 6 (c), our PointDreamer completes the missing area with high-quality texture compared to other methods.
4.6 Method Component Analysis
We here compare different implementations for sub-modules (geometry extraction, inpainting, and unprojection module) of our method by replacing one module at a time while keeping others unchanged. All experiments in this subsection are conducted on the full test set of the chair category of ShapeNetCoreV2 dataset with Gaussian noise (standard deviation = 0.005).
4.6.1 Geometry Extraction
In addition to POCO, we further explore extracting geometry by SPR and depth inpainting to investigate how geometry quality influences the final reconstructed textured mesh. Please refer to our supplementary file for more details on depth inpainting and quantitative results. Figure 7 (a) shows that noisy geometry leads to noisy texture, and our adopted POCO performs the best.
4.6.2 2D Inpainting
We compare different inpainting modules in our pipeline, including nearest interpolation, linear interpolation, DiffPIR [53], and our adopted DDNM, where DiffPIR is another diffusion-based 2D image restoration method with inpainting ability. We provide the visual comparisons in Figure 7 (b) and quantitative results in our supplementary file. As can be seen, the other three methods, though perform overall reasonably, fail to produce as clear textures as DDNM, indicating the importance of a strong inpainting module to the reconstruction performance.
4.6.3 Unprojection
We present the comparison results of different unprojection strategies in Figure  7 (c) and Table 4.4.
  • Opt. Scratch: optimize the texture atlas from scratch;
  • Naive: select the best views by direction priority;
  • NBF: select the best views by our proposed Non-Border-First strategy;
  • Opt. Naive: further optimize Naive’s obtained atlas;
  • Opt. NBF: further optimizes NBF’s obtained atlas while considering only loss in non-border areas.
Figure 7 (c) shows that the naive unprojection yields the above-mentioned artifacts from border areas, which can be slightly reduced by further optimization. Only our NBF and Opt. NBF nearly eliminate such artifacts. Table 4.4 also shows that our proposed NBF and Opt. NBF outperform other unprojection strategies.
Our NBF strategy can also work for other methods [51, 37] that reconstruct mesh textures from multiview images. For example, InstantMesh [51] reconstructs a textured 3D mesh from a single image, by first generating multiview images, and then predicting the mesh by a large reconstruction model. As mentioned, predicting colors by a network trained with 3D data can result in blurry textures, as shown in Figure 8 (a). While unprojecting the multiview images to the mesh can achieve clearer textures, as shown in Figure 8 (b), artifacts often occur at border areas, such as the border of the eye as marked in red. Our NBF strategy effectively solved such artifacts, as shown in (c). More examples are provided in (d-g).
5 Conclusion, Limitations and Future Work
Conclusion. We propose PointDreamer, a novel framework for textured mesh reconstruction from colored point cloud with SoTA performance. By utilizing diffusion-based 2D inpainting, it (1) reconstructs clear and high-quality mesh textures, addressing the common blurring issue; (2) shows high robustness against sparse, noisy, or even incomplete input, and (3) works in a zero-shot manner, requiring no extra training. We also propose a novel “Non-Border-First” strategy to unproject the colors of predicted 2D images back to 3D space. This strategy is the first to address the border-area artifact issue, which is less-explored but commonly-occurred in methods that generate 3D textures from multiview images.
Limitations and Future Work. (1) Like all other methods producing meshes from multi-view images, our method cannot perfectly color the small unseen areas. (2) The adopted inpainting module DDNM, though with SoTA performance, cannot perfectly inpaint all cases, especially those with extremely complex, irregular patterns or very fine details. In the future, we plan to develop an adaptive camera placement strategy to minimize invisible regions and better cover the entire mesh. Additionally, we aim to further investigate a relighting-supportive reconstruction approach: disentangling illumination effects from the input points’ colors, and reconstructing albedo colors and material information for the mesh, to better support relighting in graphics pipelines.
Appendix A Geometry Evaluation Results
To compare the reconstructed geometry quality of PointDreamer and baseline methods while ignoring textures, we report geometry evaluation results in Table F. Same as our manuscript, we use ShapeNetCoreV2 [44] (Chair, Car, Motorbike categories), Google Scanned Objects [45] and Omniobject3d [46] datasets. We follow POCO [4] to use commonly-used metrics of Chamfer L1-distance ×100 (CD), normal consistency (NC), and F-Score with threshold value 1% (FS). See POCO for more details about these metrics. Note that, as mentioned in our manuscript, we use POCO as the geometry extraction module for both our PointDreamer and the Texture Field baseline.
Appendix B More Visual Results
Figure 10 shows more visual comparisons of our PointDreamer against baseline methods. Figure 11 shows more visual results of our PointDreamer, including projected sparse multi-view images and inpainted dense multi-view images. We hope this can help the readers better understand our pipeline.
Appendix C Method Component Analysis
C.1 Geometry Extraction Module Replacement Experiment
To investigate how the geometry extraction module influences the final reconstructed textured mesh, we explore three different geometry extraction methods: POCO [4] (as adopted in our manuscript), SPR [7] and depth inpainting.
C.1.1 Depth-Inpainting for geometry extraction
@@
With 2D inpainting adopted as the key for our texture reconstruction, we can also introduce it to geometry reconstruction, by inpainting 2D depth maps instead of RGB images. Regarding this, we conduct experiments of depth inpainting, which refers to inpainting projected sparse 2D depth maps instead of RGB images, and then reconstructing geometry from the inpainted dense depth maps by depth fusion [54]. Specifically, we follow the following steps to reconstruct untextured meshes from input point clouds, see Figure 12:
  1. 1.
@@
  4. 4.
    We produce an untextured mesh by depth fusion based on the inpainted dense depth maps, and conduct mesh simplification [55] and Taubin Smooth [56] to it as post-processing, to get the final untextured mesh.
C.1.2 Results
We present the visual comparisons in our manuscript and quantitative results in Table F, respectively. As can be seen, POCO, as a state-of-the-art deep-learning-based surface reconstruction approach, outperforms the other two methods by a significant margin, especially regarding FID. Both SPR and Depth Inpainting suffer from noisy geometry and thus low-quality textures. This indicates that a higher-performance geometry extraction module contributes to a more refined reconstructed textured mesh.
C.2 2D Inpainting Module Replacement Experiment
We provide the quantitative results of replacing our 2D inpainting module from DDNM [24] to nearest interpolation, linear interpolation, and DiffPIR [53] in Table F. DiffPIR is another 2D-diffusion-based image restoration approach with inpainting ability. As can be seen, DDNM outperforms the other three methods thanks to its strong inpainting ability, which is consistent with the visual comparisons from Figure 4 (b) in our manuscript.
Appendix D More Experimental Results
D.1 Effect of Different K Values (Number of Viewpoints for Projection and Inpainting)
To investigate the effect of different numbers of viewpoints (denoted as $K$) for projection and inpainting, we conduct experiments on the motorbike category or ShapeNetCoreV2 dataset by using different $K$ values including 6, 8, and 20. Table F shows the distribution of cameras for each setting, together with the quantitative results. We can see that more views contribute to a slightly higher reconstruction quality. Visual comparisons in Figure 13 also show that, an insufficient number of views would produce artifacts in invisible or occluded areas, thus impacting the performance.
@@ D.1 Effect of Different K Values (Number of Viewpoints for Projection and Inpainting)
D.2 Impact of Degraded Input Quality: More Visual Comparisons
D.2.1 Anti-Noise Ability Analysis
Figure 14 presents the visual comparisons of our method and baseline methods’ reconstructed meshes from noisy or clean input point clouds. Note that the reconstructed untextured meshes of noisy and clean inputs are the same, this is because our adopted POCO was trained with noisy input, so before reconstructing geometry, we manually add noise to the clean input. Overall, we have the following observations, which are consistent with the quantitative results (Table 2) in our manuscript:
  1. 1.
  2. 2.
  3. 3.
D.2.2 Sparsity Test
Appendix E Implementation Details of Texture Field Baseline
E.1 Network architecture
Inspired by works [10, 28] that represent 3D information by a feature tri-plane, we adopt the network architecture of the open-source Convolutional Occupancy Network [40], which also follows a tri-plane representation. Specifically, we modify its encoder from taking three-dimensional input (xyz) to six-dimensional input (xyzrgb). Then, we modify its one-dimensional output head for occupancy prediction to three-dimensional for RGB color prediction.
E.2 Training
We follow 3DGen [10] to train the network by the per-3D-point MSE loss on predicted and GT colors of sampled 3D points. Additionally, we also tried to employ the per-2D-pixel MSE loss of 2D images rendered from GT meshes and predicted meshes by differentiable rendering, but the experimental results are worse, as shown in Figure 16. Therefore, we adopt the per-3D-point MSE loss in our manuscript. We train the texture field network on a single NVIDIA GeForce GTX 3090 GPU with a batch size of 24 and a learning rate of 0.0001 for 2592, 024 iterations (about 356 epochs).
E.3 Inference.
During inference, we first use POCO [4] to predict an untextured mesh from the input point cloud, and apply UV mapping to it by Xatlas [43], which produces the 3D positions of each valid pixel in the texture atlas. We query the color of each of these 3D positions by our trained texture field network, to inpaint the texture atlas, so as to obtain the final textured mesh.
Appendix F Applications for large scenes
PointDreamer, as a zero-shot method, can be adapted to indoor or outdoor scenes. Figure 9 shows two examples from Tanks and Temples Dataset [57] and MultiScan [58] dataset. Since our adopted diffusion model only supports images at $256\times 256$ resolution, we downsample the scene point cloud before reconstruction. Also, our adopted POCO is relatively slow, especially with large-scale input point clouds. In the future, we may seek to develop a more large-scale-friendly version of PointDreamer, by adopting higher-resolution diffusion models and geometry reconstruction models designed for big scenes. 2D inpainting will remain our key idea, with only submodules replaced.
TABLE V: Geometry evaluation results of our adopted POCO and baseline methods. Note that baseline Texture Field also uses POCO for geometry extraction.